The Teenager Company makes and sells skateboards at an average price of $70 each. During
the past year, they sold 4,000 of these skateboards. The company believes that the price
elasticity for this product is about −2.5. If it decreases the price to $63, what should be the
quantity sold? Will revenue increase? Why?
1
Expert's answer
2018-03-08T09:32:06-0500
P1 = $70, Q1 = 4,000 skateboards, Ed = −2.5. If P2 = $63, then the
quantity sold is:
Ed = (Q2 - Q1)/(P2 - P1)*(P2 + P1)/(Q2 + Q1),
-2.5 = (Q2 - 4,000)/(63 - 70)*(63 + 70)/(Q2 + 4,000),
-2.5 = (Q2 - 4,000)/(-7)*133/(Q2 + 4,000),
2.5 = (Q2 - 4,000)*19/(Q2 + 4,000),
5/38 = (Q2 - 4,000)/(Q2 + 4,000),
5(Q2 + 4,000) = 38(Q2 - 4,000),
33Q2 = 172,000,
Q2 = 5,212 skateboards.
The revenue will increase, because demand is elastic.